OpenStudy (vortish):

simplify the expression assume all variable are positive (k^1/6)^-15/(k^7)^1/2

OpenStudy (vortish):

\[\frac{ (k^1/6)^-15 }{(k^7)^1/2 ? }\]

OpenStudy (anonymous):

\[\huge \dfrac{ \dfrac{ (\dfrac{ k }{ 6 })^{-15} }{ k^7 } }{ 2 }\] \[\huge \dfrac{ \dfrac{ \dfrac{ 6^{15} }{ k^{15} } }{ k^7 } }{ 2 }\] \[\huge =\dfrac{ 6^{15} }{ k^{15} }\times \frac{ 2 }{ k^7 }\]